The price of a 10-year bond can decline by approximately 9% if interest rates rise by 1% point. To illustrate this, suppose you own a 10-year U.S. Treasury bond that has a bond rate of 2% per year. How much money will you lose if the value of the bond today is $10,000 (face value of the bond) and the yield increases to 3% within the next few months?
1
Expert's answer
2021-10-06T04:07:31-0400


To have less than one left, one needs a DF < 1/1000. no. 12%

This happens after 10 half-lives because (1/2)10 = 1/1024, so after 200 minutes (3 hrs 20 min).



• Ratio of total radioisotope activity to total mass of the element present

• Measured in Bq/g or Bq/mole

• Useful if radioactive and non-radioactive isotopes of the same element are in mixture (sample with carrier)

• Maximum Specific Activity in carrier-free samples:

Carrier Free Specific Activity (CFSA)


"CFSA [Bq/g] ≈ 4.8\times1018 /(A\times T1/2) for T1/2 \space given\space in\space days,"